OpenStudy (anonymous):

Please help simplify this: ((5/sqrt11) - (4sqrt33))/ 1 +(5/sqrt11)(4sqrt11)

OpenStudy (anonymous):

\[(5/\sqrt{11} - 4\sqrt{3})\div 1 + (5/\sqrt{11})\times(4\sqrt{3})\]

OpenStudy (anonymous):